OpenStudy (anonymous):

What value of b makes the statement true? 17 = b – (–6.2)

OpenStudy (anonymous):

explain and medal!

OpenStudy (jango_in_dtown):

17 = b – (–6.2) implies 17=b+12 implies b=17-12=5 so b=5 satisfies the equation

OpenStudy (jango_in_dtown):

@nward.gmail123

OpenStudy (anonymous):

Choose the answer. 8. (3 pt) What value of b makes the statement true? 17 = b – (–6.2) A. –23.2 B. –10.8 C. 10.8 D. 23.2

OpenStudy (anonymous):

it's c

OpenStudy (jango_in_dtown):

oh i see it 6.2 not 6*2=12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (jango_in_dtown):

17 = b – (–6.2) add -6.2 on both sides then you have b=17+(-6.2)=17-6.2=11.8

OpenStudy (jango_in_dtown):

10.8

OpenStudy (jango_in_dtown):

ha ha ha

OpenStudy (anonymous):

\[10.8+6.2=\]

OpenStudy (jango_in_dtown):

17

OpenStudy (anonymous):

hmm why did i get 17.6?

OpenStudy (anonymous):

ooh nevermind jiji

OpenStudy (jango_in_dtown):

because you did wrong

OpenStudy (jango_in_dtown):

:)

OpenStudy (anonymous):

i was substracting

OpenStudy (anonymous):

9. (3 pt) On a number line, point S has coordinate –5.4 and point T has coordinate 2.5. Which expressions represent the distance between point S and point T? Choose all answers that are correct. A. B. C. D.

OpenStudy (jango_in_dtown):

see a+b=c implies a=c-b

OpenStudy (anonymous):

help with a few more?plzzz

OpenStudy (anonymous):

ooopsss

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

9. (3 pt) On a number line, point S has coordinate –5.4 and point T has coordinate 2.5. Which expressions represent the distance between point S and point T? Choose all answers that are correct. A. B. C. D.

OpenStudy (anonymous):

ahhhhh i can't copy

OpenStudy (anonymous):

Choose the answer. 10. (3 pt) Sam had a full 1 L bottle of olive oil. He poured 0.25 L of the olive oil into a container to make salad dressing. He also poured an unknown amount x onto a plate as a bread dip. There was 0.6 L of olive oil remaining in the bottle. Which equation can be used to solve for x? A. 1 + 0.25 – x = 0.6 B. 1 + 0.25 + x = 0.6 C. 1 – 0.25 – x = 0.6 D. 1 – 0.25 + x = 0.6

OpenStudy (jango_in_dtown):

see the formula is |b-a| where b is the coordinate of T and a is the coordinate of s. So the required distance is |2.5-(-5.4)|=|2.5+5.4|=7.9

OpenStudy (anonymous):

that'a another one

OpenStudy (anonymous):

oooh thanks

OpenStudy (anonymous):

3 pt) Sam had a full 1 L bottle of olive oil. He poured 0.25 L of the olive oil into a container to make salad dressing. He also poured an unknown amount x onto a plate as a bread dip. There was 0.6 L of olive oil remaining in the bottle. Which equation can be used to solve for x? A. 1 + 0.25 – x = 0.6 B. 1 + 0.25 + x = 0.6 C. 1 – 0.25 – x = 0.6 D. 1 – 0.25 + x = 0.6

OpenStudy (jango_in_dtown):

see initially Sam has 1 ltr then he poured .25ltr so then amount=1-0.25 again he poured x ltr so amount remaining=1-0.25-x then he had left=0.6ltr so the equation is 1-0.25-x=0.6

OpenStudy (anonymous):

ooh i see thanks

OpenStudy (anonymous):

A!

OpenStudy (jango_in_dtown):

Its C

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Question Navigator Choose all answers that are correct. 3. (3 pt) Point C is graphed on a number line at –7. Point D is 13 units away from point C on the number line. What could be the coordinates of point D? Choose all answers that are correct. A. –20 B. 3 C. 6 D. 20

OpenStudy (anonymous):